### v2

# Imaging of electrically controlled van der Waals layer stacking in 1-TaS2

Abstract
Van der Waals (vdW) materials exhibit a variety of states that can be switched with low power at low temperatures, offering a viable cryogenic āflash memoryā required for the classical control electronics for solid-state quantum information processing. In 1-TaS2, a non-volatile metallic āhiddenā state can be induced from an insulating equilibrium charge-density wave ground state using either optical or electrical pulses. Given that conventional memristors form localized, filamentary channels which support the current, a key question for design concerns the geometry of the conduction region in highly energy-efficient 1-TaS2 devices. Here, we report in operando micro-beam X-ray diffraction, fluorescence, and concurrent transport measurements, allowing us to spatially image the non-thermal hidden state induced by electrical switching of 1-TaS2. Our results reveal a long-range ordered, non-filamentary switched state that extends well below the electrodes, implying that the self-organized, collective growth of the hidden phase is driven by a combination of charge flow and lattice strain. Our unique combination of techniques showcases the potential of non-destructive, three-dimensional X-ray imaging to study bulk switching properties in microscopic detail, namely electrical control of the vdW layer stacking.
Switching between low and high resistance states via electrical excitation has significant potential for analog and neuromorphic computing applicationsĀ [1, 2, 3]. A key requirement for this is the ability to reversibly transition between two resistance states which can occur through mechanisms such as the formation of a conductive filament, ferroelectric and magnetic tunneling junctions, or phase transitions between crystalline and amorphous states [3, 4, 5, 6]. There is substantial research primarily focused on identifying new candidates for fast and energy-efficient memory devices, and also cryo-computingĀ [7]. Van der WaalsĀ (vdW) materials, due to their layered structure, are well-suited not just to explore novel states and phase transitionsĀ [11, 12, 13, 14] but are potentially also useful for scalable electronic devicesĀ [8, 9, 10].
Among vdW materials, the transition metal dichalcogenide 1-TaS2 attracts considerable attention for its unique combination of correlated electron phenomena. These include various charge-density wave (CDW) statesĀ [15, 16], superconductivity under pressure or dopingĀ [17, 18], and a putative quantum spin liquid phaseĀ [19, 20]. At low temperatures it is believed to exhibit a Mott insulating stateĀ [21, 17], characterized by commensurateĀ (C) CDW order formed by star-shaped polaron domains that tessellate the layers below Ā KĀ [22, 21]. Moreover, and of greatest interest to us here, it shows non-thermal, reversible switching to a metastable, metallic āhiddenāĀ (H) CDW phase upon application of ultrashort optical or electrical pulses. The lifetime of this HCDW state is short at elevated temperatures and stable below Ā KĀ [23, 24, 25, 26, 27].
The microscopic origin of the non-thermal switching of 1-TaS2 has been intensely studied but is not fully understoodĀ [23, 28, 29, 24, 30, 31, 32, 33]. It can be described in terms of a low-temperature free-energy landscape that features a global minimum (the Mott insulating equilibrium state) and multiple local minima separated by potential energy barriersĀ [23, 24, 33, 32]. Excitation with a laser or current pulse can drive the system into these metastable states.
Optical switching of 1-TaS2 with femtosecond laser pulses, exciting electron and holes homogeneously, has been extensively studied with both quasi-static [30, 34, 35] and time-resolvedĀ [36, 23, 37, 38, 31, 39, 40, 41, 42] techniques. Surface-sensitive scanning tunneling microscopyĀ (STM) shows formation of conducting domain wallsĀ [23, 30], while bulk-sensitive X-ray diffraction (XRD) reveals changes in the out-of-plane polaron stacking and a different in-plane commensurability of the HCDWĀ state compared to the low- and room-temperature CDW states, confirming a non-thermal switching mechanismĀ [31, 42].
In contrast, electrical switching has been less explored despite its promise for application in highly-efficient cryo-memory devices due to the compatibility with standard electronics. The electrically-induced hidden (e-HCDW) state has been studied using transport measurementsĀ [26, 25, 43, 44], STMĀ [27, 29, 28], and angle-resolved photoemission spectroscopyĀ [45], revealing macroscopic behavior similar to the optically-induced hidden (o-HCDW) state. Domain wall formation at the surface has also been observed with STM, but unlike the o-HCDW, it is understood that the e-HCDW state results from excitation due to charge carrier separation, leading to more inhomogeneous domain wallsĀ [27, 32]. Furthermore, in contrast to the o-HCDW state, the e-HCDW can be triggered with much longer pulses up to Ā [46], suggesting key differences in their switching mechanisms. As previously discussedĀ [27, 32], the switching is non-thermal even for long current pulses.
It is unclear whether optical and electrical excitations lead to the same local free-energy minimum. If they do, can these two distinct perturbationsādiffering in pulse duration, electron-hole pairs accessed, and directionalityāfollow the same non-thermal pathway, or are there different mechanisms at play? Also, the lack of spatially-resolved studies on the e-HCDW leaves the question unanswered whether the switching is surface-localized or propagates through the bulk. Likewise, it is unknown if the e-HCDW spreads across the entire region in-between the electrodes or forms conducting filament channels. Finally, it is not established whether the principal actor of the switching is the electric field or the current. Resolving the spatial structure and population of the e-HCDW would provide insights into the switching mechanisms, and thereby contribute to designing devices with improved performance and scalability.
The required information can be obtained by mapping the bulk CDW states before and after switching. If the e-HCDW is structurally equivalent to the o-HCDW, then the known CDW diffraction peaks from photo-driven experimentsĀ [31] can be used as a fingerprint to identify the e-HCDW switching region. Here we report on in situ resistance measurements with micro-beam X-ray fluorescenceĀ (XRF) and diffractionĀ (XRD) to spatially image a nano-fabricated 1-TaS2 device in operando (Fig.Ā 1). Analysis of CDW Bragg peaks reveals the detailed three-dimensionalĀ (3D) spatial evolution of the material properties before and after electrical excitation.
Results
FigureĀ 1b shows an optical image of the 1-TaS2 device, formed by a contacted flake, used to investigate the e-HCDW state (fabrication details in Methods). The distance between the electrodes was set toĀ m, considering the m-sized X-ray spot and that the switching voltage scales with the distance between the electrodesĀ [43]. There are two inert contacts between the outer electrodes that are not used for the resistance measurement. The device is located using XRF (Fig.Ā 1c) which is sensitive to even small traces of embedded elementsĀ [47, 48] (Supplementary Information). Before and after applying current pulses, XRD and XRF measurements are taken at room temperature and , as well as different energies spanning an extended portion of reciprocal space (Fig.Ā 1d,e). Conversion to 3D reciprocal space allows us to identify the equilibrium and newly appearing signals, as well as to determine their position and peak shapes (Supplementary Information). Temperature-dependent resistance of the device is measured in a two-probe configuration (Fig.Ā 1f). Upon cooling, the flake goes through a first-order phase transition, resulting in a step-like increase of the resistance at . We denote the unswitched low-temperature CCDW state with A. To induce the e-HCDW state, single square current pulses with a width of are applied. The pulse amplitude is increased gradually until a resistance drop occurs. We denote the state B with resistance compared to A as partially-switched, reached by applying a pulse amplitude ofĀ . Increasing the pulse height toĀ induces another resistance drop to of A, which we refer to as fully-switchedĀ C. Following the resistance upon heating, we observe the characteristic relaxation from the HCDW state to the equilibrium high-temperature states, as well as the first-order, hysteretic transitionĀ [24], confirming that the HCDW state is induced in the device. As a function of time, the device resistance exhibits changes upon cooling, heating, and current pulse application, else it is stable over hours. During switching, we observe some resistance increase, likely due to a partial relaxation to the CCDW state, and then, as the current amplitude is further increased, a resistance drop indicating that the HCDW state is induced. We note that devices with an optimized design can be switched with a single current pulse [43, 44], whereas with our device design, optimized for in situ XRD measurements, multiple pulses with increasing amplitude have to be applied. During the XRD measurements at a constant temperature of the resistance remains stable. Importantly, the resistances measured at room temperature before and after the switching and XRD measurements are identical, which directly proves the non-destructive nature of our technique.
Several structural changes occur in the non-equilibrium o-HCDW state compared to the equilibrium CCDW order: (i) the CCDW and ādimerā peaks vanishĀ [31] (the latter are associated with inter-layer dimerization of the star-like domains), (ii) a new long-range order (at different positions in reciprocal space) appears as HCDW peaksĀ [31, 30], and (iii) the out-of-plane lattice constant contractsĀ [49] (observed as a shift of the lattice peaks in the out-of-plane direction). In the following, we show that these three characteristic features are also observed for the e-HCDW state. Here, we focus on the HCDW signal, whereas the vanishing dimer CCDW peaks and the out-of-plane lattice contraction are addressed in the Supplementary Information.
For each position on the device, we reconstruct the 3D reciprocal space and look for a decrease in intensity of the CCDW peak, as well as newly appearing signals hinting at the e-HCDW state: three 3D regions of interest (ROI) around the (013) lattice, as well as nearby CCDW and o-HCDW positionsĀ [31] are set, and the intensity within these ROIs is integrated. The shape of the flake can be clearly seen in the lattice peak maps (Fig.Ā 2a,d,g), which remains intact throughout the switching process. However, a change occurs in the CDW structure of the flake: the CCDW peak map before switching (A, Fig.Ā 2b) shows the same outline as the lattice peak; but Fig.Ā 2e,h in the partially- and fully-switched states, B and C, respectively, reveal that the CCDW peak intensity in the bottom left part of the flake vanishes. Concurrently, an e-HCDW peak signal appears in the same region in-between the electrodes (Fig.Ā 2f,i). That is, not the entire flake switches from the CCDW to the HCDW order, but rather a conducting channel appears between the electrodes on one edge of the flake, while the rest of the device remains in the equilibrium state.
Having identified the spatial region of the flake that switches to the HCDW state, we examine the momentum-space structure and real-space distribution of the CDW states. The CCDW signal in Fig.Ā 3a features the characteristic elongation parallel to the out-of-planeĀ direction due to partial disorder, and the reciprocal space position is close to the previously reported values [50, 51, 31] (Tab.Ā 1). FiguresĀ 3b-d are from the fully-switched stateĀ C, deduced from different regions on the flake and revealing that CCDW order gradually gives way to the e-HCDW state as the switching channel is approached. The reciprocal space position of the e-HCDW peak in-between the electrodes is also found to be consistent with the o-HCDW signal in the literatureĀ [31] (Tab.Ā 1). Therefore, we conclude that the optically- and electrically-driven HCDW order are equivalent not only from an electronic but also a 3D structural point of view. In turn, this also implies control of the switching into the same local minimum of the free-energy landscape regardless of the excitation method.
Next, we use the intensity shift from the CCDW to the HCDW peak to assess the switching depth. We take the ratio of the HCDW and the total (HCDW CCDW) signals as a proxy for the volume fraction of the switched layers. Since the Pd/Au electrodes are on top of the flakeāthus, the highest current density is thereāwe also assume that the switching starts from the top. FigureĀ 4Ā shows the respective 3DĀ tomographic representation of the device. Clearly, the switching is not restricted to the surface but penetrates deep into the bulk of the 500-nm-thick flake. This observation is in agreement with previous reports on the importance of the out-of-plane layer reconfiguration in the HCDW stateĀ [29, 31, 42]. We take cuts parallel and perpendicular to the electrode gap to further characterize the bulk material switching: The HCDW region extends in volume going from the partially- B to the fully-switched stateĀ C. Parallel to the gap the cut reveals that the switching starts at the edge of the flake that is the shortest path to ground. Moreover, from the cut perpendicular to the gap we see that the electrons flowing from the to the electrode do not symmetrically switch the intergap space. In addition, a fraction of remains unswitched between the electrodes. Though using XRD, we cannot extract information down to the single layer and consequently the remaining unswitched intergap layers are within the error of our technique. We also observe sizeable HCDW order induced under the electrodes that is not directly the shortest path to ground, hinting that not only the out-of-plane layer reorderingĀ [29, 31, 42] but also strain and the out-of-plane conductivityĀ [46, 52] play an important role in the switching process.
Discussion
We non-destructively acquire in operando ātomogramsā of cryo-memory device switching, addressing a long-standing challenge for the engineering of phase-change memory devicesĀ [3, 53]. Our approach allows us to identify the switched HCDW state along one edge of the flake and extending under the electrodes into the bulk of the material. This hints at charge injection as the initiator of the switching. The appearance of a current-induced conducting boundary channel at the flake edge / electrode interface aligns also with previous findings [54].
Finite element method simulations (Supplementary Information) indicate that the geometry of the device dictates the switching channelās location: The switching occurs where the electrode gap is narrowest, providing the shortest path to ground and resulting in the highest current density during pulse application. Once the channel has formed, the current primarily flows through it, and the switched volume only extends when the current is further increased, i.e., from state B to C (Fig.Ā 4). The AuĀ electrodes ensure high conductivity along the flake, rendering a conductivity gradient unlikelyĀ [55]. The switching channel also does not seem to form along defects or impurities, as its size ( ) exceeds that of defects or domainsĀ [56, 30]. Moreover, the lattice peak maps in Fig.Ā 2a,d,g confirm that the flake remains intact during the switching process, ruling out structural fracture as the root cause.
The switching region can also be located by mapping the vanishing dimer peak (Supplementary Information), associated with the breaking of the pairwise alignment of vdW layers [31]. The nature of the equilibrium CCDW stateāwhether it is a Mott or a band insulatorāremains controversial and is found to be tied to the out-of-plane restacking [42]. Here, we confirm control and local modification of the out-of-plane stacking via current pulses.
In- and out-of-plane strain between the flake and the substrate or electrodes may not only influence the stability of the HCDW stateĀ [57, 24] but also the switching regionās location. Studies on other phase-change materials show that local strain inhomogeneities can arise upon electrical triggeringĀ [58]. We also observe an out-of-plane lattice contraction localized exclusively in the intergap space of our 1-TaS2 device when a current pulse is applied (Supplementary Information). This creates local strain between the switched and unswitched areas. Thus, we identify strain propagation via the CDW domains as an essential factor contributing to the localized and bulk-nature of the switching, in close analogy to what occurs for ferroelectric devicesĀ [59]. Since strain is a long-range effect, it may also explain why the switching extends beneath the electrodesāan aspect not captured by the strain-free simulations.
Interestingly, the original CCDW intensity is higher in the region where the switching channel forms. Observed only for the CCDW but not the lattice peak (Fig.Ā 2a,b), this indicates that structural differences, such as variations in flake thickness, are not responsible but rather that switching is potentially promoted by a particularly coherent (well-ordered) CDW state. Our results, therefore, underscore the importance of resolving electronic heterogeneities and looking into the role they play in the non-thermal switching process of 1-TaS2.
Our 3D reconstruction of the phase change suggests optimized device designs using narrow flakes with electrodes positioned along the edges. Ultimately, we envision separate electrode pairs that allow for multiple switched regions and controlled cross talk between them, eventually enabling highly-efficient and fast logic operations.
Studying bulk switching using XRD with nano/micro-beams offers many opportunities. Direct imaging of filamentary paths arising from strain or different structural reconfiguration can also be applied to other memristive materials. So far these have mostly been investigated in thin films without contacts [60], in-plane with transmission electron microscopy, or indirectly and destructively using conductive atomic force microtomographyĀ [61, 62, 63, 64]. Our non-destructive, 3D, in operando approach represents a significant advance over these technologies, providing microscopic detail of the switching mechanism, thereby advancing the design of scalable next-generation memory technologies.
Methods
Device preparation.
1-TaS2 single-crystals were synthesized using chemical vapor transport with iodine as the transport agent leading to consistent crystal qualityĀ [19].
Flakes were mechanically exfoliated from bulk crystals using GelPak and excess flakes removed using Scotch tape such that a large, single flake could be deposited onto an oxidized Si wafer (oxide thickness:Ā ) with predefined Ti/AuĀ alignment markers. The electrodes were written on the flake using electron beam lithography, followed by physical vapor deposition of Ā Pd and Ā Au. The investigated flake has lateral dimensions of about 50 and a thickness of . The in-plane crystal orientation of the device was characterized at the Material Science beamline of Swiss Light Source synchrotron [65], and it was then glued onto a CuĀ sample holder using GEĀ varnish.
Transport measurements and sample cooling.
The device was mounted in a CryoVac 4HeĀ cryostat with Kapton windows for the incident and scattered X-rays. The resistance was measured in a two-probe configuration on the flake, and converted to four-probes on the sample holder such that only the resistance of the Pd/Au electrodes on the flake is measured in series with the flake. The electrical setup consisted of a Keithley 6221 pulsed current source and a Keithley 2182 nano-voltmeter connected in delta mode. For electrical switching single, square-wave pulses were applied using the current source. The sample temperature was monitored using a Cernox thermometer mounted on the sample holder about from the device. A specific cooling procedure ensured that the first-order phase transition to the CCDW state was crossed in a controlled fashion. The following ramp rates were used: 300 to :Ā /, 250 to : /, 200Ā toĀ : /, and 100 to : /.
X-ray diffraction and fluorescence. Spatially-resolved X-ray scattering was performed at the microXAS beamline of the Swiss Light SourceĀ [66], allowing for simultaneous in situ resistance measurements, as well as XRD and XRF imaging. The incidence angle of the X-rays with respect to the device was fixed to , resulting in a beam spot size (v h) using Kirkpatrick-Baez focusing mirrors. The nominal X-ray flux was adjusted to about Ā photons/s. Considering that the X-rays were passing through about air before impinging on the sample, the effective incident flux on the sample was about and lower at 9.1 and , respectively. As the cryostat was mounted on translation stages without rotational degree of freedom, to take XRD spatial maps we scanned through reciprocal space by changing the incoming X-rays energy between nd using the undulator and monochromator. The diffraction and fluorescence signals were recorded by Eiger X 4M and VIAMP-KC detectors, respectively. To locate the device the XRF signal at the Au () and Ta () -edges was used (Supplementary Information).
Data availability
Data supporting the figures and other findings of this study are available from the corresponding authors upon request.
Acknowledgements.
The authors acknowledge the Paul Scherrer Institute, Villigen, Switzerland for provision of synchrotron radiation at the microXAS and Material Science beamlines of the Swiss Light Source. We thank P.Ā Sutar for synthesizing the 1-TaS2 crystals and acknowledge technical support from the PSIĀ PICO operations team. Furthermore, we are grateful to B.Ā Meyer, M.Ā Birri and S.Ā Stutz for technical support before and during the beamtimes, as well as fruitful discussions with N.Ā Taufertshƶfer. This research was funded by the Swiss National Science FoundationĀ (SNSF) and the Slovenian Research And Innovation AgencyĀ (ARIS) as a part of the WEAVE framework Grant NumberĀ 213148 (ARIS project N1-0290). R.V., A.M. and D.M. thank ARIS for funding the research program P1-0400, and D.M. thanks ARIS for funding the research program J7-3146. B.L. and M.T. thank ARIS for funding the research program P2-0415. C.B., W.H., A.G.M. and G.A. acknowledge funding from the European Research Council under the European Unionās Horizon 2020 research and innovation program, within Grant Agreement 810451 (HERO). N.H. received funding from the European Unionās Horizon 2020 research and innovation program under the Marie Sklodowska-Curie Grant Agreement 884104 (PSI-FELLOW-III-3i).Author contributions
C.B., D.M. and S.G. conceived the project with input from G.A. and Y.E. C.B. and D.K. fabricated the device with input from D.S. C.B., A.M., N.H., S.-W.H., F.D. and S.G. characterized the device. C.B., N.H., D.F.S., R.V., F.D., D.G. and S.G. prepared the experiment. C.B., N.H., D.F.S., W.H., H.G.B., R.V., F.D., S.-W.H., A.G.M., and S.G. carried out the experiment with input from D.G. C.B., W.H., H.G.B., and N.H. analyzed the data. B.L. conducted the simulations with input from M.T., C.B., A.M., D.M. and S.G. C.B., N.H., D.M. and S.G. wrote the manuscript with input from all co-authors.
Competing interests
The authors declare no competing interest.
| Peak | Literature | Experiment | Exp. uncertainty |
| Lattice | |||
| CCDW | |||
| HCDW |
Supplementary information
Imaging of electrically controlled van der Waals layer stacking in 1-TaS2
Corinna Burri
Nelson Hua Dario Ferreira Sanchez
Wenxiang Hu HenryĀ G.Ā Bell
RokĀ Venturini
Shih-WenĀ Huang
Aidan G. McConnell FarisĀ DizdareviÄ
Anže Mraz Damjan Svetin
Benjamin LipovÅ”ek Marko TopiÄ
DimitriosĀ Kazazis
GabrielĀ Aeppli
Daniel Grolimund
Yasin Ekinci
Dragan Mihailovic
Simon Gerber
January 29, 2025
I Fluorescence
X-ray fluorescence of Ta and Au was used to locate the device and set the range of the spatially-resolved scans. The signal was captured by VIAMP-KC detectors. Even small traces of elements such as Cu from the sample holder, Ti from the binary markers on the substrate, and Ni traces under the electrodes were detected (Fig.Ā 5). Fluorescence data were fitted using the PyMCA librariesĀ [67].
II X-ray diffraction data processing
II.1 Experimental geometry constraint and reciprocal space reconstruction
In the following, we describe how we assigned pixels on the detector at each X-ray energy (, , E) to reciprocal space coordinates (h, k, l).
The known parameters in this experiment are the energy of the X-rays, the orientation of the detector which was perpendicular to the X-ray beam and further detector specifications of the Eiger X 4M detector. In addition, we know the optimal pixel and energy of scattering reflections of three accessible lattice peaks (004), (013), (014) and four commensurate CCDW peaks on the detector and also their theoretical (h, k, l) parameters [51], leading to seven known quantities. The measured peak positions were determined by fitting a Lorentzian to the projected intensities. The CCDW peaks are elongated along the out-of-plane direction, therefore only the in-plane (h, k) dimensions were fitted. The unknown parameters in this setup were the relative position of the detector with respect to the sample (x,Ā y,Ā z) which was roughly measured and used as an initial guess, the sample orientation parameterized with angles () and the lattice constants of the flake (a, b, c). If we parametrize the Eiger X 4M orientation with Cartesian coordinates (x,Ā y,Ā z) where (x, y) point along the x, y pixel directions and z perpendicular to the detector surface, the flake orientation can be expressed in this coordinate system using angles (). denotes the angle between the surface normal of the detector and the surface normal of the flake, is the azimuthal angle and the tilt angle. Since there were only seven measured parameters (three lattice and four CCDW reflections), we took the in-plane lattice parameters as known [69] and only the out-of-plane lattice parameter c as unknown.
As a next step, a cost function was defined that includes all known and unknown parameters, and with which the full experimental geometry was described as
where are the unknown parameters, are the known parameters and the two-norm is used.
With this function, the reciprocal space coordinates of our measured peaks was compared to the theoretically known position [51, 31]. The cost function was minimized with respect to using a standard minimization algorithm from the SciPy optimization python library [70] which finds the experimental conditions of the unknown parameters that matches the experimental results. The resulting error table is shown in Tab.Ā 2 with small errors, meaning that there is good agreement between the measured peaks and the known structure of the material.
| Peak | Literature (r.l.u) | Experiment (r.l.u) | Two-norm error (pixels) |
| (013) CCDW1 | 1.5 | ||
| (013) CCDW2 | 0.7 | ||
| (014) CCDW1 | 1.2 | ||
| (014) CCDW2 | 0.7 | ||
| (013) lattice | 3.5 | ||
| (004) lattice | 5.0 | ||
| (014) lattice | 2.3 |
Once this experimental geometry was calculated, the real space position with respect to the sample of all pixels on the detector were defined. The reciprocal space coordinates (h, k, l) were assigned to the pixels on the detector at every X-ray energy using the Laue equation. We note that the pixel size of the Eiger X 4M detector is whereas the lateral size of the flake was about . Hence, any movement of the diffraction pattern due to rastering over the sample was still within 1 pixel of the detector and can be neglected.
The conversion from (, , E) to (h, k, l) leads to a curved surface in reciprocal space which was probed by the flat 2D detector in real space. Putting all the curved surfaces of the measured energies together, allowed us to reconstruct the 3D shape of the lattice, CCDW, HCDW and dimer peaks. FigureĀ 6 shows this 3D reciprocal space representation measured in the fully-switched stateĀ C. This is a volumetric plot using the intensities of the measured peaks, giving an overview over the peak shapes and position. Importantly, this is not a spatially-resolved 3DĀ reciprocal image, it is the reconstruction coming from the sum of all pixels measured at the different energies, therefore we observe all peaks in the same plot. Next to the overview plot are extracts of reciprocal space around the lattice and CCDW peak in the unswitchedĀ A and in the fully-switched stateĀ C. We see that in the fully-switched state a HCDW peak appeared below the CCDW peak, as well as an additional lattice satellite (Sect.Ā V). For the spatially-resolved intensity maps of the device, we used such 3D reconstruction of every pixel. The peak positions were determined from the 3DĀ reconstruction by summing along two reciprocal space directions to obtain a 1D intensity plot along the remaining reciprocal space direction and fitting a Lorentzian function to the peak.
II.2 Gold ring correction
We observed five ring-like features in the raw diffraction patterns (Fig.Ā 7a). After the detector position with respect to the sample was determined, the angle of these rings, i.e. the radius , at a certain energy of the incoming X-rays was determined using Braggās law
where is the wavelength of the incoming X-rays, Ā the spacing between the crystal planes, an integer for the diffraction order and the wavenumber. Comparison with literature showed that the radii of the rings matches Au with Miller indices (111), (200), (220), (311) and (222) [71] (Fig.Ā 7c). The electrodes are made from amorphous gold, therefore we observe these rings when spatially scanning over the device. Taking a region of interest (ROI) with radius ] (in m) corresponding to the thickness of the gold rings, allowed us to collect and sum up the intensity of all detector pixels within this ROI. We did this procedure for every pixel ( and ) on the sample which gave us spatially-resolved intensity maps outlining the electrodes of the device. We used these electrode images for aligning the spatial scans at the different X-ray energies.
After confirming that no peak of interest cut through the gold rings, we replaced them for the 3D reciprocal space reconstruction by an average background (Fig.Ā 7b). This was simply done for better visibility of the peak shapes in the reconstruction, but did not affect the reconstruction itself.
II.3 Spatially-resolved maps
In the following, we describe how we obtain the spatially-resolved intensity maps, as well as the spatially-resolved projections in reciprocal space. For each state of the device, i.e. the room-temperature, as well as the unswitched A, partially- B and fully-switched statesĀ C at the base temperature of 6Ā K, we measured spatial maps at different energies. As a next step, we performed a spatial jitter correction. That is, we spatially aligned the scans at different energies and for the different states. We did this by choosing a gold ring which shows clear outlines of the electrodes in the spatial maps. For each state, we first chose one reference spatial mapĀ measured at a certain X-ray energy. We then aligned all other spatial maps measured at different energies to this reference map. Due to the finite scan range, we confined a possible shift in the x and y spatial directions to and (in pixel) and found such that
summing over the pixel indices , was minimized. The result shows in general good alignment within ±1 pixel. But there were a few scans at certain energies which still had to be manually tweaked to improve the alignment.
After having the scans at different energies for every state spatially aligned and a 3D reciprocal space reconstruction for each pixel performed, we defined 3D ROIs in reciprocal space around the peaks of interest. Since the CCDW and dimer peaks are elongated along l, as has been reported before [31], we took cylindrical ROIs around these peaks. For the lattice and HCDW peaks which are better defined along the out-of-plane direction, we took spherical ROIs. We computed the signal within these ROIs by summing up the intensities. For every spatial map at every energy, we set a spatial 1.5 (room temperature) and () grid according to the minimal step size of the scan. Then, we assigned intensity values to the grid by considering that if the closest scan to a respective grid pixel is 2Ā pixel, we set it to zero and otherwise we consider the data within the predefined ROI in reciprocal space. For the spatially-resolved intensity maps, such as in Fig.Ā 2 of the main text, the intensity in the spatial pixels was therefore the sum of the intensities within the ROI. We note that with a minimal step size in the and direction of at room temperature and at the base temperature, we were in fact oversampling given the X-ray spot size of 1.5 (v h).
After identifying the region which switched to the HCDW state from the spatial maps, we could separate the device into different spatial regions to extract the spatially-resolved 2D reciprocal space projections (Fig.Ā 3 of the main text) and calculate the ratio of the HCDW and the CCDW peak intensity. The spatial region near the intergap space (light purple) corresponds to a intensity ratio and that region in the switching channel (dark purple) corresponds to a ratio of . Then, we summed up the 3D reciprocal space reconstructions of the pixels within these regions. For the projected maps, such as Fig.Ā 3 of the main text, we set a grid in reciprocal space and interpolate. From this, we obtain the 2D projections by summing along one reciprocal space axis and dividing by the number of pixels in the spatial regions.
III Room-temperature spatial maps
At the beginning of the experiments, the device was mounted at room temperature and spatial scans at different X-ray energies were measured to check the alignment, as well as to record the lattice and NCCDW signals. We took spatially-resolved scans at a few energies, corresponding to only selected cuts through reciprocal space. Hence, we only observed a few peaks with reduced statistics, compared to the low-temperature measurements. This resulted in a less accurate 3D reconstruction. Nevertheless, we took 3D ROIs around the (013) lattice peak and a respective NCCDW peak and plotted the corresponding spatial intensity map which both reproduce the general shape of the flake (Fig.Ā 8).
IV Dimers and phase switching
As discussed in [31, 72], dimer peaks associated with a doubling of the real-space unit cell in the out-of-plane direction were observed in the CCDW state and vanish upon switching to the o-HCDW state [31]. We also observed dimer peaks in our measurements (Fig.Ā 6) and took 3D ROIs around those in reciprocal space. We recorded only a few energy scans cutting through the dimer peaks, meaning that the statistics was reduced compared to other peaks. FigureĀ 9 shows the spatially-resolved intensity maps of the CCDW and dimer peak in the unswitched A, the partially- B and the fully-switched state C. We see that both peak intensities vanish in the bottom left part of the flake in the fully- switched state. Hence, we observe the same vanishing dimer peaks in the e-HCDW state in the intergap space as reported for the o-HCDW state, meaning for both there is also collapse of dimerization. Indirectly, observed as a local doubling of the out-of-plane dispersion, this has also been reported in a recent ARPES study [45]. As addressed in the main text, we observed that the intensity of the CCDW and dimer peaks in the unswitched state was the highest in the bottom left part of the flake that switches to the HCDW state. This supports the notion that the location of the switching channel may not only be dictated by the geometry of the device, but also the most homogeneous (dimerized) CDW order.
V Switching-induced out-of-plane lattice contraction
As reported before for the o-HCDW state [49], we also observed a change in the out-of-plane lattice constant in the e-HCDW state. FigureĀ 10 shows the 1D projection onto the out-of-plane reciprocal space coordinate l of the (004) lattice peak. The signal was measured on the entire flake (no spatial resolution) in the unswitchedĀ A, the partially- B and the fully-switched state C. The high intensity peak is the (004) lattice peak reflection. We note that because this measurement is not spatially-resolved, intensity (from the unswitched portions of the flake) still shows up there even in the C state. Clearly, in the B and C states a satellite lattice peak emerges from the HCDW state in the switching channel. Thus, we observed that for the e-HCDW state the out-of-plane lattice peak position shifts to a higher reciprocal space value, meaning that the lattice contracts by . The out-of-plane lattice constant difference between the switched and unswitched region leads to strain along the border which we speculate facilitates the switching in-plane but also in volume. Similarly, the c lattice constant contracts between the NCCDW and CCDW states [73, 74, 75, 72].
We set two ROIs around the lattice signals in reciprocal space: and r.l.u. for the main and satellite peak, respectively, to obtain spatial maps by summing up the intensity within these ROIs (Fig.Ā 11). Also here we observe that the HCDW state switching channel appeared at the bottom left corner of the flake. Thus, not only the electronic CDW structure is modified in the e-HCDW state but also the out-of-plane lattice contracts in the non-volatile switching channel of the device. In the following, we show the agreement of the HCDW state switching channel observed from the out-of-plane lattice contraction and the HCDW peak.
As done for Fig.Ā 3 of the main text, we define different spatial regions of the device: outside of the switching channel (light blue), near the switching channel (light purple) and in-between the electrodes (dark purple). We then map out the 2D reciprocal space projections of the (004) lattice peak (Fig.Ā 12). Unlike the CCDW peak which shows a characteristic elongation, the lattice peak is sharp also in the out-of-plane direction. The correlation of the ratio of the projected out-of-plane lattice intensities of the main and satellite signal, , and that of the HCDW and CCDW peak intensity, , is shown in Fig.Ā 13a. FigureĀ 13b compares the contours of the switching region based on the ratio of the lattice signals, and the HCDW/CCDW peaks, revealing that the two are also spatially correlated. Thus, also the lattice peak contraction can be used as a fingerprint to locate the HCDW switching.
VI Finite element method simulations
Numerical simulations were performed using the COMSOL Multiphysics software, which uses the finite element method to solve partial differential equations governing heat conduction in solid materials and devices:
where is the temperature, the volume and the heat source term. The density of the material , heat capacity [76], as well as the thermal [77] and electrical conductivity of all materials in the device were taken from literature or experiments [23, 32]. The thermal model was coupled to the electromagnetic one which provided in the differential equations by calculating the resistive losses in the device, i.e. Joule heating.
The structure and dimensions of the simulated device is shown in Fig.Ā 14a. The geometry of the Au contacts and the 1-TaS2 crystal flake was realistically replicated both in shape and size, including irregularities such as the narrower spacing at one (left) side of the structure and the two additional disconnected contact fingers in the middle. Excitation of the device was applied via the two outer contacts by assuming square wave current pulses of amplitudes between and . At the beginning of each simulation, the device was fully in the unswitched state A at temperature. After applying the excitation, switching of the elements from A to the fully-switched state C was handled by a simplified model which predicts formation of the HCDW state domain wall network under the influence of a sufficiently large charge carrier injection. Thus, in the present simulations the state of each element was switched once the current density of the element reached a predefined current density threshold of , calibrated by the experiment. FigureĀ 14b shows the switched material (orange) for different magnitudes of current excitation. All cases identify the preferential area of device switching at the left side of the device, which is influenced primarily by the narrower spacing between the contacts. With increasing current amplitude, the area gradually expands along the intergap space. Furthermore, we observe that the penetration depth of the switching is determined by the level of anisotropy of 1-TaS2. In the isotropic case (Fig.Ā 14b, top row), the switched volume propagates throughout the entire thickness of the crystal flake, whereas it remains closer to the surface when assuming an out/in-plane conductivity anisotropy of 1:5 (Fig.Ā 14b, middle row). A final set of simulations (Fig.Ā 14b, bottom row), we tested also the influence of the disconnected contact fingers. While they do affect the quantitative extent of the switched volume, the additional conductive paths they provide qualitative agree with the simulations in the top row of Fig.Ā 14b.
Contrary to the experiment these simulations do not reproduce the widening of the switched volume laterally and beneath the contacts. We speculate that this effect in the experiment is driven by propagation of local strain between the switched and unswitched areaāaĀ feature presently not included in the numerical simulations. We also observe that compared to the experiment, the device switches fully at a lower current amplitude (between to ). We attribute this to slight discrepancies between the assumed and effective material parameters. Finally, in Fig.Ā 14c, device heating due to the current excitation is also demonstrated for the isotropic situation without influence of the disconnected middle contacts (Fig.Ā 14b, top row). The uneven heating upon current pulse application localized to the left side of the device is also in agreement with the preferential path of the current flow across the 1-TaS2 flake.